A tank can be filled by a pipe in 3 hours and emptied by another pipe in 4 hours. How much time will be required to fill an empty take if both are running?
Answer by richwmiller(17219)

1/a+1/b=1/x
1/3+1/-4=1/x
-4/-12+3/-12=1/x
-1/-12=1/x
x=-12/-1
x=12 hours working together
check
x/a+x/b=1
x/3+x/-4=1
12/3+12/-4=1
4+-3=1
1=1
ok